OpenStudy (anonymous):

for x not equal to zero what choice of f(0) will make the functions continuous at x=0 1) f(x) = [(1+x)^1/3-1]/x 2) f(x) = [5|x|-3 tan x + sin 2x]/x

OpenStudy (amistre64):

take the limit as x approaches zero from the left and right ....

OpenStudy (anonymous):

would you plz show a few steps ? @amistre64

OpenStudy (amistre64):

trying to think of a simple way to approach it ...

OpenStudy (amistre64):

can we use taylor series?

OpenStudy (amistre64):

squeeze thrm maybe?

OpenStudy (anonymous):

have no idea about taylor series, not taught

OpenStudy (amistre64):

well, brute math says we table it .... approach it from the left and right and log the values

OpenStudy (amistre64):

but that requires a lot of calculation

OpenStudy (amistre64):

\[\sqrt[3]{1+x} = 1+\frac x3-\frac{x^2}{9}+...\] \[\sqrt[3]{1+x}-1 = \frac x3-\frac{x^2}{9}+...\] \[\frac{\sqrt[3]{1+x}-1}{x} = \frac 13-\frac{x}{9}+...\] the right side zeroes out and we are left with 1/3

OpenStudy (amistre64):

but you say you dont know that method

OpenStudy (anonymous):

the value for f(0) for the first problem is given as 1/3 for the second it will never be continuous !

OpenStudy (amistre64):

yeah the second one, by graph is not a hole, but a jump; so yeah nothing could fill it in

OpenStudy (anonymous):

ok... i understand what you are trying to explain !! so when we have x = 0 all other terms except 1/3 will be zero

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

for the second one suppose i write statements ..what should be the statements to conclude that will never be continuous

OpenStudy (amistre64):

well, sin(u)/u = 1 as u to 0 5|x|-3 tan x + sin 2x ------------------ x 5|x| 3 tan x sin 2x ---- - ------- + ------- x x x 5|x| 3 sinx sin 2x 2 ---- - ------- + ------- * -- x cosx x x 2 5|x| 3 ---- - ----- + 2 x cosx 5|x| ---- - 3 + 2 x 5|x| ---- - 1 as x to 0 x

OpenStudy (anonymous):

how did 3/cos x become 3 ? did not understand that part !!

OpenStudy (amistre64):

so, as x to 0, 5x/x - 1 = 4 -5x/x - 1 = -6 cos(0) = 1

OpenStudy (amistre64):

since we have the limit approaching 4 and -6 at the same time, there is no point that will be able to bridge the gap

OpenStudy (anonymous):

thanks a lot ..you are a hero :)

OpenStudy (amistre64):

youre welcome ... the first one i dont see any other way about it :/

OpenStudy (anonymous):

but i understood the first one , it was like binomial expansion .. thanks again !!

OpenStudy (amistre64):

binomial expansion yes :) good luck

OpenStudy (anonymous):

i am closing the question